Portfolio Average Return Standard Deviation Beta
A 14.7% 18.6% 1.47
B 8.8 14.2 .78
C 11.2 16.0 1.22

The risk-free rate is 4.5 percent and the market risk premium is 7 percent.

What is the Treynor ratio of a portfolio comprised of 30 percent portfolio A, 20 percent portfolio B, and 50 percent portfolio C?
Answer #1

Expected Return of Portfolio = 0.30(0.147) + 0.20(0.088) + 0.50(0.112)

Expected Return of Portfolio = 11.77%

Beta of Portfolio = 0.30(1.47) + 0.20(0.78) + 0.50(1.22)

Beta of Portfolio = 1.207

Treynor Ratio = (0.1177 - 0.045)/1.207

Treynor Ratio = 0.06